# A distributed augmented Lagrangian decomposition algorithm for constrained optimization

### v2

@@ Abstract
Within the framework of the Augmented Lagrangian (AL), we introduce a novel distributed optimization method called Distributed Augmented Lagrangian Decomposition (DALD). We provide a rigorous convergence proof for the standard version of this method, which is designed to tackle general constrained optimization problems. To address the high iteration costs in early stages, we propose several accelerated variants of DALD that enhances efficiency without compromising theoretical guarantees, supported by a comprehensive convergence analysis. To facilitate the description of the distributed optimization process, the concept of hierarchical coordination networks is introduced, integrating hierarchical matrix concepts to aid in this explanation. We further explore and expand the applicability of the DALD method and demonstrate how it unifies existing distributed optimization theories within the AL framework. The effectiveness and applicability of the proposed distributed optimization method and its variants are further validated through numerical experiments.
Index Terms:
augmented Lagrangian, decomposition, distributed optimization, multi-agent systems
@@ I Introduction
However, the aforementioned distributed optimization algorithms typically rely on varying degrees of assumptions, particularly regarding the separability and convexity of the objective function, the existence or linearity of constraints, and are often limited to a two-level hierarchy in the sequence of subproblem solving. These assumptions are fundamental to their theoretical foundations and practical performance. Moreover, these algorithms typically adhere to a fixed subproblem-solving sequence, either simply serial or parallel, without delving deeply into the internal structure of the problem. Against this backdrop, this paper introduces a novel distributed optimization technique, the Distributed Augmented Lagrangian Decomposition (DALD) method, designed for distributed solutions to more general optimization problems. What sets this method apart is its ability to handle objective functions that may include non-separable terms, without requiring linear constraints. During the development of this method, we also reveal the essence of existing distributed optimization algorithms based on the AL techniques.
@@ II-A Related Work
$\underset{\begin{subarray}{c}\mathbf{x}\in X\end{subarray}}{\text{min}}$ $f(\mathbf{x})=\sum_{i=1}^{n}f_{i}(\mathbf{x}_{i})+f_{0}(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n})$ (1)
 s.t. $\psi(\mathbf{x})=\psi(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n})\leq 0,$
where the functions $f:\mathbb{R}^{N}\rightarrow\mathbb{R}$, $\psi:\mathbb{R}^{N}\rightarrow\mathbb{R}^{m}$ are continuously differentiable and convex. Furthermore, $\mathbf{x}_{i}$ are multidimensional subvectors of $\mathbf{x}$, and $\mathbf{x}_{i}=[x_{i1},x_{i2},\ldots,x_{iN_{i}}]^{\top}\in X_{i}\subseteq\mathbb{R}^{N_{i}}$, $i\in\mathbb{N}_{n}=\{1,2,\ldots,n\}$. Collectively, $\mathbf{x}=[\mathbf{x}_{1}^{\top},\mathbf{x}_{2}^{\top},\ldots,\mathbf{x}_{n}^{\top}]^{\top}\in X\subseteq\mathbb{R}^{N}$ with $N=\sum_{i=1}^{n}N_{i}$, and $X$ is a Cartesian product of closed sets: $X=X_{1}\times X_{2}\times\cdots\times X_{n}$, where $X_{i}$ denotes a given nonempty closed subset of $\mathbb{R}^{N_{i}}$, $i\in\mathbb{N}_{n}$. In the objective function of (1), $f_{0}(\mathbf{x})$ represents the cumulative sum of all non-separable sub-terms among different subproblems.
To facilitate further analysis, we introduce a slack vector $\mathbf{e}\in\mathbb{R}^{m}$ to handle the constraints, embedding it into the modified constraint function defined as $\varphi(\mathbf{x})=\psi(\mathbf{x})+\;\mathbf{e}^{2}.$ Although the function notation $\varphi(\mathbf{x})$ explicitly depends only on $\mathbf{x}$, it implicitly incorporates the slack vector $\mathbf{e}$, which will be optimized jointly with $\mathbf{x}$. By introducing Lagrange multipliers $\mu\in\mathbb{R}^{m}$ and penalty parameters $\rho\in\mathbb{R}^{m}$, we relax problem (1) and define the Lagrangian as:
@@
The symbol $\circ$ represents the Hadamard product: $\mathbf{a}\circ\mathbf{b}=[a_{1},a_{2},\ldots,a_{n}]^{\top}\circ[b_{1},b_{2},\ldots,b_{n}]^{\top}=[a_{1}b_{1},a_{2}b_{2},\ldots,$ $a_{n}b_{n}]^{\top}$, and $\|\cdot\|$ denotes the $\ell_{2}$-norm.
@@ II-B Notation
$\min$ $\underbrace{5v_{1}+2v_{2}+v_{3}+3v_{4}}_{\sum_{i=1}^{n}f_{i}(\mathbf{x}_{i})}+\underbrace{v_{5}v_{7}+v_{6}v_{8}}_{f_{0}(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n})}$
s.t. $\varphi(\mathbf{x})=\left\{\begin{aligned} &v_{1}+v_{5}v_{6}-7=0,\\ &v_{2}+v_{5}+v_{7}-10=0,\\ &v_{5}v_{6}+e^{v_{7}}-12=0,\end{aligned}\right$.
@@
    We define $f_{0}(\mathbf{x}_{i},\mathbf{x}_{-i})$ as the sum of non-separable terms within $f_{0}(\mathbf{x})$ with respect to subproblem $i$ and its coupling subproblems, which is referred to as the coupling objective function. Let $f_{i}(\mathbf{x}_{i})$ be the local objective function for subproblem $P_{i}$, and $f_{i}(\mathbf{x}_{i},\mathbf{x}_{-i})=f_{i}(\mathbf{x}_{i})+f_{0}(\mathbf{x}_{i},\mathbf{x}_{-i})$, $i\in\mathbb{N}_{n}$. Note that $f_{0}(\mathbf{x})$ is not equivalent to $\sum f_{0}(\mathbf{x}_{i},\mathbf{x}_{-i})$. To illustrate, we have $f_{1}(\mathbf{x}_{1})=5v_{1},f_{0}(\mathbf{x}_{1},\mathbf{x}_{-1})=v_{5}v_{7},f_{i}(\mathbf{x}_{1},\mathbf{x}_{-1})=5v_{1}+v_{5}v_{7},f_{0}(\mathbf{x})=v_{5}v_{7}+v_{6}v_{8}.$
@@
    Constraint $\phi_{i}(\mathbf{x}_{i},\mathbf{x}_{-i})=0$, which contains the local variable $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$, is called the coupling constraint of subproblem $P_{i}$ and subproblem $P_{j}$, $i\in\mathbb{N}_{n}$, $j\in\mathcal{R}_{i}$. As clearly indicated in 2) of Definition 1, when subproblems are coupled, their coupling constraints may share common elements. For example, $\phi_{2}(\mathbf{x}_{2},\mathbf{x}_{-2})$ consists of the elements in the set $\{v_{2}+v_{5}+v_{7}-10=0,\,v_{5}v_{6}+e^{v_{7}}-12=0\}$ and $\phi_{3}(\mathbf{x}_{3},\mathbf{x}_{-3})$ consists of the elements in the set $\{v_{1}+v_{5}v_{6}-7=0,\,v_{5}v_{6}+e^{v_{7}}-12=0\}.$
@@ III-B Convergence Analysis of Standard DALD
    The functions $f(\mathbf{x})$ and $\psi(\mathbf{x})$ are continuously differentiable and convex.
@@
    The Lagrangian (2) has a saddle point $(\mathbf{x}^{*},\mu^{*})\subseteq\mathbb{R}^{N}\times\mathbb{R}^{m}$:
@@
Assumptions (A1)(A3) imply that (4) and (12) are convex, and the strong duality holds for the problem (1), i.e., the optimal values of the primal and dual problems are equal. Under these assumptions, in particular, when all constraints $\psi(\mathbf{x})$ are affine functions and a feasible solution exists, the Slater condition is automatically satisfied, ensuring that (A3) holds. Furthermore, these assumptions guarantee the convergence of the ALM. Assumption (A4) ensures the smooth implementation of the inner loops of DALD.
@@
Taking the limit as $v\to\infty$, we get
@@
Using the optimality condition of Prop. 3.1.1 in [37], we have
@@ IV-B Parameters Determination and Accelerated Version
Lemma 1 Suppose $\mathbf{x}^{*}$ is the optimal point of problem (1). Then, there exists a constant vector $\gamma=(\gamma_{1},\gamma_{2},\ldots,\gamma_{n})\subseteq\mathbb{R}^{N}$, where $\gamma_{i}\subseteq\mathbb{R}^{N_{i}}$, such that $\nabla_{i}L(\mathbf{x}^{*},\mu^{*})=\nabla_{i}\Lambda_{\rho}(\mathbf{x}^{*},\mu^{*})=\gamma_{i},i\in\mathbb{N}_{n}$. Specifically, when $X=\mathbb{R}^{N}$, $\gamma=0$.
@@
where $\gamma_{i}\subseteq\mathbb{R}^{N_{i}},i\in\mathbb{N}_{n}$. Furthermore, (18) yields
@@
s.t. $\psi(\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n})\leq 0$,
@@ IV-C Framework Unification
For typical BCD methods, different component algorithms can be selected in the solver layer based on the characteristics of the subproblems. For instance, if the subproblems are differentiable, gradient-based algorithms such as classic Gradient Descent (GD) or Newton’s method (NM) can be considered. For example, in the case of the classic gradient descent algorithm (see Appendix VII-C for details), the gradient of the objective function with respect to the elements $x_{ij}$ in block $\mathbf{x}_{i}$ needs to be computed, involving all information related to $x_{ij}$, $i\in\mathbb{N}_{1}$, $j\in\mathbb{N}_{N_{i}}$. From the perspective of decomposition and coordination, gradient descent algorithms can also be viewed as a generalized distributed optimization algorithm [9]. According to (28), first-order optimization algorithms that use gradient information (such as GD) are implemented in the solver layer in a manner consistent with Jacobian-type standards.
@@
For clarity, Table I provides a comparison to highlight the differences among the involved algorithms. It can be observed that DALD handles a more general range of problem forms and is more versatile in terms of its termination criteria. It is important to note that the algorithms shown in Fig. 4 are mature classical algorithms proposed by previous researchers, and their variants are not depicted in the figure. We believe that the problem forms and coordination mechanisms addressed by these variants are essentially similar to those of the prototype algorithms and fall within the interpretive scope of DALD. For instance, most variants of the ADMM are still confined to problems with fully separable and convex objectives, linear constraints, or 2-level coordination.
@@ V-B Network Flow Problem
Consider a directed graph $G=(\mathcal{N},A)$, where $\mathcal{N}$ represents the vertices and $A$ the edges. Each edge $(i,j)$ has an associated cost $c_{ij}$, and the flow through this edge is denoted by $t_{ij}$. The flow $t_{ij}$ must satisfy constraints $a_{ij}\leq t_{ij}\leq b_{ij}$. The supply at node $i$ is given by $s_{i}$, where $a_{ij}$, $b_{ij}$, and $s_{i}$ are predefined. For source nodes $S$, we have $s_{i}>0$ for $\forall i\in S$, while for sink nodes $D$, $s_{i}<0$, $\forall i\in D$. Consider the problem:
@@ VI Conclusion
This paper investigates a distributed solver-agnostic optimization technique, DALD, which is distinguished by its ability to solve optimization problems with general characteristics. Specifically, we introduce both the standard and accelerated versions of the method, providing the corresponding convergence proofs. In addition, we indirectly address a longstanding issue in the field of distributed optimization: why the direct extension of ADMM is not necessarily convergent. Unlike previous research, we introduce the concept of “solver layer” and highlight that DALD can leverage different solvers to solve subproblems without depending on any specific solver. We also briefly review and analyze existing distributed optimization methods, revealing that they fall within the explanatory scope of the DALD framework. Different from Professor Bertsekas’ perspective [37, Ch. 7.4.2], we view DALD as a distributed implementation of alternating optimization within the framework of the AL for constrained optimization problems.

### v3

@@ Abstract
Within the framework of the augmented Lagrangian (AL), we propose a novel distributed optimization method, termed Distributed Augmented Lagrangian Decomposition (DALD), and provide a rigorous convergence proof for its standard version. To address the high iteration costs in early stages, we propose several accelerated variants of DALD that enhances efficiency without compromising theoretical guarantees, supported by a comprehensive convergence analysis. To facilitate the description of the distributed optimization process, the concept of hierarchical coordination networks is introduced, integrating hierarchical matrix concepts to aid in this explanation. We further explore and expand the applicability of the DALD method and demonstrate how it unifies existing distributed optimization theories within the AL framework. The effectiveness and applicability of the proposed distributed optimization method and its variants are further validated through numerical experiments.
@@ I Introduction
However, the aforementioned distributed optimization algorithms typically rely on varying degrees of assumptions, particularly regarding the separability and convexity of the objective function, and are often limited to a two-level hierarchy in the sequence of subproblem solving. These assumptions are fundamental to their theoretical foundations and practical performance. Moreover, these algorithms typically adhere to a fixed subproblem-solving sequence, either simply serial or parallel, without delving deeply into the internal structure of the problem. Against this backdrop, this paper introduces a novel distributed optimization technique, the Distributed Augmented Lagrangian Decomposition (DALD) method, designed for distributed solutions to more general optimization problems. What sets this method apart is its ability to handle objective functions that may include non-separable terms, while enabling distributed optimization over flexible communication topologies. During the development of this method, we also reveal the essence of existing distributed optimization algorithms based on the AL techniques.
@@ II-A Related Work
$\underset{\begin{subarray}{c}\mathbf{x}\in X\end{subarray}}{\text{min}}$ $f(\mathbf{x})$ (1)
 s.t. $\sum_{i=1}^{n}\mathbf{A}_{i}\mathbf{x}_{i}=\mathbf{b}$
where the function $f:\mathbb{R}^{N}\rightarrow\mathbb{R}$ is continuously differentiable and convex. Furthermore, $\mathbf{x}_{i}$ are multidimensional subvectors of $\mathbf{x}$, and $\mathbf{x}_{i}=[x_{i1},x_{i2},\ldots,x_{iN_{i}}]^{\top}\in X_{i}\subseteq\mathbb{R}^{N_{i}}$, $i\in\mathbb{N}_{n}=\{1,2,\ldots,n\}$. Collectively, $\mathbf{x}=[\mathbf{x}_{1}^{\top},\mathbf{x}_{2}^{\top},\ldots,\mathbf{x}_{n}^{\top}]^{\top}\in X\subseteq\mathbb{R}^{N}$ with $N=\sum_{i=1}^{n}N_{i}$, and $X$ is a Cartesian product of closed sets: $X=X_{1}\times X_{2}\times\cdots\times X_{n}$, where $X_{i}$ denotes a given nonempty closed subset of $\mathbb{R}^{N_{i}}$, and each $\mathbf{A}_{i}$ is a matrix of dimension $m\times N_{i}$, $\mathbf{b}\in\mathbb{R}^{m}$ is a vector, $i\in\mathbb{N}_{n}$.
To facilitate the symbolic representation for subsequent analysis, the constraints in problem (1) can be formulated as:
$\varphi(\mathbf{x})=\sum_{i=1}^{n}\mathbf{A}_{i}\mathbf{x}_{i}-\mathbf{b}=0,$
where $\varphi:\mathbb{R}^{N}\rightarrow\mathbb{R}^{m}$. Subsequently, by introducing Lagrange multipliers $\mu\in\mathbb{R}^{m}$ and penalty parameters $\rho\in\mathbb{R}^{m}$, we relax problem (1) and define the Lagrangian as:
@@
The symbol $\circ$ represents the Hadamard product: $\mathbf{c}\circ\mathbf{d}=[c_{1},c_{2},\ldots,c_{n}]^{\top}\circ[d_{1},d_{2},\ldots,d_{n}]^{\top}=[c_{1}d_{1},c_{2}d_{2},\ldots,c_{n}d_{n}]^{\top}$, and $\|\cdot\|$ denotes the $\ell_{2}$-norm.
@@ II-B Notation
$\min$ $\underbrace{v_{3}+v_{4}+v_{5}^{2}+v_{7}^{2}+v_{6}^{2}+v_{8}^{2}}_{\sum_{i=1}^{n}f_{i}(\mathbf{x}_{i})}\;+\;\underbrace{2v_{5}v_{7}+2v_{6}v_{8}}_{f_{0}(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n})}$
s.t. $\varphi(\mathbf{x})=\left\{\begin{aligned} &v_{1}+v_{5}+v_{6}-7=0,\\ &v_{2}+v_{5}+v_{7}-10=0,\\ &v_{5}+v_{6}+2v_{7}-12=0,\end{aligned}\right$.
@@
    In the objective function, we define $f_{0}(\mathbf{x})$ as the cumulative sum of all non-separable sub-terms across different subproblems. We further define $f_{0}(\mathbf{x}_{i},\mathbf{x}_{-i})$ as the sum of the non-separable terms in $f_{0}(\mathbf{x})$ that are associated with subproblem $i$ and its coupling subproblems, which is referred to as the coupling objective function. Let $f_{i}(\mathbf{x}_{i})$ be the local objective function for subproblem $P_{i}$, and $f_{i}(\mathbf{x}_{i},\mathbf{x}_{-i})=f_{i}(\mathbf{x}_{i})+f_{0}(\mathbf{x}_{i},\mathbf{x}_{-i})$, $i\in\mathbb{N}_{n}$. Note that $f_{0}(\mathbf{x})$ is not equivalent to $\sum f_{0}(\mathbf{x}_{i},\mathbf{x}_{-i})$. To illustrate, we have $f_{1}(\mathbf{x}_{1})=v_{5}^{2},f_{0}(\mathbf{x}_{1},\mathbf{x}_{-1})=2v_{5}v_{7},f_{1}(\mathbf{x}_{1},\mathbf{x}_{-1})=v_{5}^{2}+2v_{5}v_{7},f_{0}(\mathbf{x})=2v_{5}v_{7}+2v_{6}v_{8}.$
@@
    Constraint $\phi_{i}(\mathbf{x}_{i},\mathbf{x}_{-i})=0$, which contains the local variable $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$, is called the coupling constraint of subproblem $P_{i}$ and subproblem $P_{j}$, $i\in\mathbb{N}_{n}$, $j\in\mathcal{R}_{i}$. As clearly indicated in 2) of Definition 1, when subproblems are coupled, their coupling constraints may share common elements. For example, $\phi_{2}(\mathbf{x}_{2},\mathbf{x}_{-2})$ consists of the elements in the set $\{v_{2}+v_{5}+v_{7}-10=0,\,v_{5}+v_{6}+2v_{7}-12=0\}$ and $\phi_{3}(\mathbf{x}_{3},\mathbf{x}_{-3})$ consists of the elements in the set $\{v_{1}+v_{5}+v_{6}-7=0,\,v_{5}+v_{6}+2v_{7}-12=0\}.$
@@ III-B Convergence Analysis of Standard DALD
    The function $f(\mathbf{x})$ is continuously differentiable and convex.
@@
    The Lagrangian (2) has a saddle point $(\mathbf{x}^{*},\mu^{*})\in\mathbb{R}^{N}\times\mathbb{R}^{m}$:
@@
Assumptions (A1)(A3) imply that (4) and (12) are convex, and the strong duality holds for the problem (1), i.e., the optimal values of the primal and dual problems are equal. Under these assumptions, since all constraints $\varphi(\mathbf{x})$ are affine functions, and when a feasible solution exists, the Slater condition is automatically satisfied, thereby guaranteeing that (A3) holds. Furthermore, these assumptions guarantee the convergence of the ALM. Assumption (A4) ensures the smooth implementation of the inner loops of DALD.
@@
Taking the limit as $v\to\infty$, we obtain
@@
Since $\mathcal{F}$ is convex, its restriction to any coordinate (with the other variables fixed) is also convex. By the first-order optimality condition (Prop. 3.1.1 in [37]), it follows that
@@
Remark 1: For problem (1), suppose that $\varphi(\mathbf{x})$ represents general constraints (not limited to affine functions). Then either the AL in (3) remains convex, in which case Theorem 1 still holds, or every limit point of $\{\mathbf{x}^{k,v}\}$ is a critical point of $\Lambda_{\rho}(\mathbf{x},\mu^{k})$ over $X$.
@@ IV-B Parameters Determination and Accelerated Version
Lemma 1 Suppose $\mathbf{x}^{*}$ is the optimal point of problem (1). Then, there exists a constant vector $\gamma=(\gamma_{1},\gamma_{2},\ldots,\gamma_{n})\in\mathbb{R}^{N}$, where $\gamma_{i}\in\mathbb{R}^{N_{i}}$, such that $\nabla_{i}L(\mathbf{x}^{*},\mu^{*})=\nabla_{i}\Lambda_{\rho}(\mathbf{x}^{*},\mu^{*})=\gamma_{i},i\in\mathbb{N}_{n}$. Specifically, when $X=\mathbb{R}^{N}$, $\gamma=0$.
@@
where $\gamma_{i}\in\mathbb{R}^{N_{i}},i\in\mathbb{N}_{n}$. Furthermore, (18) yields
@@
s.t. $\sum_{i=1}^{n}\mathbf{A}_{i}\mathbf{x}_{i}=\mathbf{b}$,
@@ IV-C Framework Unification
For typical BCD methods, different component algorithms can be selected in the solver layer based on the characteristics of the subproblems. For instance, if the subproblems are differentiable, gradient-based algorithms such as classic Gradient Descent (GD) or Newton’s method (NM) can be considered. For example, in the case of the classic GD (see Appendix VII-C for details), the gradient of the objective function with respect to the elements $x_{ij}$ in block $\mathbf{x}_{i}$ needs to be computed, involving all information related to $x_{ij}$, $i\in\mathbb{N}_{1}$, $j\in\mathbb{N}_{N_{i}}$. From the perspective of decomposition and coordination, gradient descent algorithms can also be viewed as a generalized distributed optimization algorithm [9]. According to (28), first-order optimization algorithms that use gradient information (such as GD) are implemented in the solver layer in a manner consistent with Jacobian-type standards.
@@
For clarity, Table I provides a comparison to highlight the differences among the involved algorithms. It can be observed that DALD handles a more general range of problem forms and is more versatile in terms of its termination criteria. It is important to note that the algorithms shown in Fig. 4 are mature classical algorithms proposed by previous researchers, and their variants are not depicted in the figure. We believe that the problem forms and coordination mechanisms addressed by these variants are essentially similar to those of the prototype algorithms and fall within the interpretive scope of DALD. For instance, most variants of the ADMM are still confined to problems with fully separable and convex objectives, or 2-level coordination.
@@ V-B Network Flow Problem
Consider a directed graph $G=(\mathcal{N},A)$, where $\mathcal{N}$ is the set of vertices and $A$ is the set of directed edges. Each edge $(i,j)$ has an associated cost $c_{ij}$, and the flow through this edge is denoted by $t_{ij}$. The flow $t_{ij}$ must satisfy constraints $a_{ij}\leq t_{ij}\leq b_{ij}$. The supply at node $i$ is given by $s_{i}$, where $a_{ij}$, $b_{ij}$, and $s_{i}$ are predefined. For source nodes $S$, we have $s_{i}>0$ for $\forall i\in S$, while for sink nodes $D$, $s_{i}<0$, $\forall i\in D$. Consider the problem:
@@ VI Conclusion
This paper investigates a distributed, solver-agnostic optimization technique, DALD, which is distinguished by its ability to solve optimization problems with broader characteristics under more flexible communication topologies. Specifically, we introduce both the standard and accelerated versions of the method and provide the corresponding convergence proofs. In addition, we indirectly address a longstanding issue in the field of distributed optimization: why the direct extension of ADMM is not necessarily convergent. Unlike previous research, we introduce the concept of “solver layer” and highlight that DALD can leverage different solvers to solve subproblems without depending on any specific solver. We also briefly review and analyze existing distributed optimization methods, revealing that they fall within the explanatory scope of the DALD framework. Different from Professor Bertsekas’ perspective [37, Ch. 7.4.2], we view DALD as a distributed implementation of alternating optimization within the framework of the AL for constrained optimization problems.